### v1

@@ 2 Introduction
Here, we demonstrate quantum-coherent motional coupling and the direct observation of a topological Berry phase in a fully tunable 2D microtrap array consisting of three ions arranged in an equilateral triangle (Fig. 1). We prepare all nine modes of motion close to their ground states and inject a single phonon to create an entangled Bell state of the coupled eigenmodes. By dynamically tuning the couplings between motional modes, we can continuously deform the character of the single-phonon motional state to encircle a conical intersection in the motional eigenvalue landscape, evidenced by a topological Berry phase [berry_phase_1984] visible in single-phonon interference experiments. To the best of our knowledge, Berry phases from conical intersections have previously only been inferred from the observation of node-lines in wavepacket distributions in quantum simulations [wang_2023, valahu_2023, whitlow_2023].
@@ 3 The Triangular Ion Microtrap Array
where $m$ is the mass of a 9Be+ ion. The rate of change of the scaled curvatures with the applied potentials is $\delta s_{A}/\delta V_{A}=\delta s_{B}/\delta V_{B}\approx$ 1.2 mV-1. By symmetry the curvatures in sites B and C change at equal amounts $\alpha~\delta s_{A}/\delta V_{A}$ with $\alpha\approx-0.383$ and likewise for A and C when $\delta V_{B}$ is changed (also see Methods).
@@ 4 Static Interaction Tuning
We then apply a weak oscillating electric field pulse at the ion positions using the chip-integrated antenna and scan its frequency to probe the motional eigenmode frequencies. When the tone is resonant with one of the normal modes, the mode will be excited into a small coherent state [mccormick_2019_coherent] with a maximal average motional occupation of $\bar{n}_{j}\approx 1$. We detect the presence or absence of this motional excitation using a Raman red sideband pulse on both ions followed by fluorescence readout of the ion spin state. The normalized fluorescence of the two ions (which is correlated with the probability that the motion was resonantly excited by the probe electric field) is shown in Fig. 1. The bright lines trace the normal mode frequencies in good agreement with the theoretical predictions based on the eigenvalues of Eq. (1) that are shown by colored lines (see Methods for details). We observe the degeneracy of two of the modes at $\delta V_{A}=0$ mV expected from the trap symmetry and the predicted gap to the highest mode frequency. The mode eigenvector $\mathbf{c}_{1}$ remains equal to $1/\sqrt{2}(0,-1,1)$ throughout with $\delta\omega_{1}$ increasing linearly with $\delta V_{A}$, as expected. Small deviations from the theory are caused by slow drifts in the applied and stray potentials on the order of tens of microvolts that occur on the time scale of hours.
@@ 7 Acknowledgments and Statement of Work
J.F.N. and D.L. performed the experiments and analyzed the data, with N.K.L. assisting. K.C.M. and J.K. performed preliminary experiments and testing of the ion trap. C.W.H. and M.G.B. led the fabrication of the ion trap. R.S., R.J., and D.L. designed the trap’s electrode geometry. S.L.T. and D.H.S., with assistance from R.J., designed and built the cryogenic ion trap system. D.J.W., A.C.W. and D.L. advised the scientific direction of the project during its development and, with D.H.S., secured its funding. D.H.S. provided technical guidance for apparatus and experimental design. J.F.N. and D.L. wrote the manuscript with input from all authors. D.L. supervised the project and developed the theoretical models of the experiments.
@@ 8.3 Observing a Berry phase
A neighborhood of 12 data points spanning approximately $\delta f=\pm 2$ kHz around each of the observed resonance peaks in Fig. 1 for fixed $\delta V_{A}$ and varying excitation frequency $f_{R}+\delta f$ was fitted to a Gaussian that yielded the center excitation frequency difference $\delta f_{0}$ and the standard deviation of the Gaussian fit. Resonances close to the crossing ($\delta V_{A}=\{-0.5,0,0.5\}$ mV) where two of the peaks are unresolved in the data were removed, because their convolution produces a single peak with a position that does not easily yield the frequencies of the two underlying resonances. The remaining peak centers approximately form three lines along the ridges observed in Fig. 1. The model for the eigenmode frequencies is derived from $\mathbf{H}(s_{A},s_{B})$ by substituting
@@
which yields $\Delta f=(3.299\pm 0.02)$ kHz (intercept) and $\Delta f\,c=(-3.97\pm 0.02)$ kHz/mV (slope). The isolated parameters can now be determined: $c=\Delta f\,c/\Delta f=(-1.202\pm 0.009)$ 1/mV and $\alpha=\Delta f\,\alpha\,c/(\Delta fc)=(-0.383\pm 0.003)$, where the uncertainties are derived by propagating the uncertainties of the fit parameters through the expressions stated above. By symmetry, $c$ and $\alpha$ also govern the change of frequencies when applying $\delta V_{B}\neq 0$; $s_{B}=c\,\delta V_{B}$. The frequency eigenvalues as a function of $\delta V_{A}$ as determined by the fit are plotted as solid lines in Fig. 1 and Fig. 2 and are in excellent agreement with the data. The fitted function is plotted without any free parameters in Fig. 2. The same fit parameters are also used in the theoretical expressions of the bright population in Fig.s 22. The fitted phonon exchange result not shown in 22 are displayed in Fig. S1. The fitted values are also converted to curvatures according to Eq.(8.2) and used in Eq.(1) to plot the curvature surfaces in Fig. 1, the intended paths for the Berry phase experiments in Fig.s 3, 3 and the additional paths shown in the supplemental information.

### v2

@@ 2 Introduction
Here, we demonstrate quantum-coherent motional coupling and the direct observation of a topological Berry phase in a fully tunable 2D microtrap array consisting of three ions arranged in an equilateral triangle (Fig. 1). We prepare all nine modes of motion close to their ground states and inject a single phonon to create an entangled Bell state of the coupled eigenmodes. By dynamically tuning the couplings between motional modes, we can continuously deform the character of the single-phonon motional state to encircle a conical intersection in the motional eigenvalue landscape, evidenced by a topological Berry phase [berry_phase_1984] visible in single-phonon interference experiments. To the best of our knowledge, Berry phases from conical intersections have previously only been inferred from the observation of node lines in wavepacket distributions in quantum simulations [wang_2023, valahu_2023, whitlow_2023].
@@ 3 The Triangular Ion Microtrap Array
where $m$ is the mass of a 9Be+ ion. The rate of change of the scaled curvatures with the applied potentials is $\delta s_{A}/\delta V_{A}=\delta s_{B}/\delta V_{B}\approx$ 1.2 mV-1. By symmetry the curvatures in sites B and C change at equal amounts $\alpha~\delta s_{A}/\delta V_{A}$ with $\alpha\approx-0.383$ and likewise for A and C when $\delta V_{B}$ is changed (see Methods).
@@ 4 Static Interaction Tuning
We then apply a weak oscillating electric field pulse at the ion positions using the chip-integrated antenna and scan its frequency to probe the motional eigenmode frequencies. When the tone is resonant with one of the normal modes, the mode will be excited into a small coherent state [mccormick_2019_coherent] with a maximal average motional occupation of $\bar{n}_{j}\approx 1$. We detect the presence or absence of this motional excitation using a Raman red sideband pulse on both ions followed by fluorescence readout of the ion spin state. The normalized fluorescence of the two ions (which is correlated with the probability that the motion was resonantly excited by the probe electric field) is shown in Fig. 1. The bright lines trace the normal mode frequencies in good agreement with the theoretical predictions based on the eigenvalues of Eq. (1) that are shown by colored lines (see Methods for details). We observe the degeneracy of two of the modes at $\delta V_{A}=0$ mV expected from the trap symmetry and the predicted gap to the highest mode frequency. The mode eigenvector $\mathbf{c}_{1}$ remains equal to $1/\sqrt{2}(0,-1,1)$ throughout with $\delta\omega_{1}$ increasing linearly with $\delta V_{A}$, as expected. Small deviations from the theory are caused by slow drifts in the applied and stray potentials on the order of tens of microvolts
@@ 7 Acknowledgments and Statement of Work
J.F.N. and D.L. performed the experiments and analyzed the data, with N.K.L. assisting. K.C.M. and J.K. performed preliminary experiments and testing of the ion trap. C.W.H., M.G.B., and E.J.H. led the fabrication of the ion trap. R.S., R.J., and D.L. designed the trap’s electrode geometry. S.L.T. and D.H.S., with assistance from R.J., designed and built the cryogenic ion trap system. D.J.W., A.C.W. and D.L. advised the scientific direction of the project during its development and, with D.H.S., secured its funding. D.H.S. provided technical guidance for apparatus and experimental design. J.F.N. and D.L. wrote the manuscript with input from all authors. D.L. supervised the project and developed the theoretical models of the experiments.
@@ 8.3 Observing a Berry phase
A set of 12 data points spanning approximately $\delta f=\pm 2$ kHz around each of the observed resonance peaks in Fig. 1 for fixed $\delta V_{A}$ and varying excitation frequency $f_{R}+\delta f$ was fitted to a Gaussian that yielded the center excitation frequency difference $\delta f_{0}$ and the width of the Gaussian fit. Resonances close to the crossing ($\delta V_{A}=\{-0.5,0,0.5\}$ mV) where two of the peaks are unresolved in the data were removed, because their convolution produces a single peak with a position that does not easily yield the frequencies of the two underlying resonances. The remaining peak centers approximately form three lines along the light-colored regions in Fig. 1. The model for the eigenmode frequencies is derived from $\mathbf{H}(s_{A},s_{B})$ by substituting
@@
which yields $\Delta f=(3.299\pm 0.02)$ kHz (intercept) and $\Delta f\,c=(-3.97\pm 0.02)$ kHz/mV (slope). The isolated parameters can now be determined: $c=\Delta f\,c/\Delta f=(-1.202\pm 0.009)$ 1/mV and $\alpha=\Delta f\,\alpha\,c/(\Delta fc)=(-0.383\pm 0.003)$, where the uncertainties are derived by propagating the uncertainties of the fit parameters through the expressions stated above. By symmetry, $c$ and $\alpha$ also govern the change of frequencies when applying $\delta V_{B}\neq 0$; $s_{B}=c\,\delta V_{B}$. The frequency eigenvalues as a function of $\delta V_{A}$ as determined by the fit are plotted as solid lines in Fig. 1 and Fig. 2 and are in excellent agreement with the data. The fitted function is plotted without any free parameters in Fig. 2. The same fit parameters are also used in the theoretical expressions of the bright population in Fig.s 22. Additional fitted phonon exchange results not shown in 22 are displayed in Fig. S1. The fitted values are also converted to curvatures according to Eq.(8.2) and used in Eq.(1) to plot the curvature surfaces in Fig. 1, the intended paths for the Berry phase experiments in Fig.s 3, 3 and the additional paths shown in the supplemental information.